# U-shaped and Inverted-U Scaling behind Emergent Abilities of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have been shown to exhibit emergent abilities in some downstream tasks, where performance seems to stagnate at first and then improve sharply and unpredictably with scale beyond a threshold. By dividing questions in the datasets according to difficulty level by average performance, we observe U-shaped scaling for hard questions, and inverted-U scaling followed by steady improvement for easy questions. Moreover, the emergence threshold roughly coincides with the point at which performance on easy questions reverts from inverse scaling to standard scaling. Capitalizing on the observable though opposing scaling trend on easy and hard questions, we propose a simple yet effective pipeline, called Slice-and-Sandwich, to predict both the emergence threshold and model performance beyond the threshold.
@@ 1 Introduction
Large language models (LLMs) (Team et al., 2023; Achiam et al., 2023; Brown, 2020; Touvron et al., 2023a; b; Workshop et al., 2022; Li et al., 2023; Jiang et al., 2024) have shown strong potential in various downstream applications (Jumper et al., 2021; Fawzi et al., 2022; Naveed et al., 2023; Kaddour et al., 2023). Though the training-loss scaling law has been well established (Kaplan et al., 2020; Hoffmann et al., 2022), the literature is inconclusive regarding how performance on downstream tasks scales. In particular, for certain downstream tasks (Srivastava et al., 2023; Lin et al., 2022a; Pilehvar & Camacho-Collados, 2019), LLMs seem to display emergent abilities: performance is stagnant even when model size scales up hundredfold, and then exhibits sharp improvement at a seemingly unpredictable critical threshold (Wei et al., 2022; Schaeffer et al., 2024a).
@@
Prior work (Schaeffer et al., 2024a; b; Lu et al., 2024) has identified crude performance metrics as a contributing factor to LLM’s apparent emergent abilities because of its inability to capture improvements of smaller models. Hu et al. (2023) proposes the PASSUNTIL metric, according to which models slowly improve with scale instead of stagnating. Schaeffer et al. (2024a) finds that LLMs display emergent abilities mainly on string-match and multiple-choice tasks (Schaeffer et al., 2024a), for which the traditional performance measure of accuracy exhibits strong discontinuity. They propose using a continuous metric such as Brier Score (Brier, 1950) or linear metric such as token edit distance (TED) (Schaeffer et al., 2024a) to better predict LLM’s scaling behavior for downstream tasks. Schaeffer et al. (2024b) further ranks several performance metrics in correlation with the model scale.
@@
Another focus of prior literature is the predictability of model performance on traditional metrics like accuracy. Wei et al. (2022) characterizes emergent abilities as unpredictable performance soar. Some studies (Ruan et al., 2024; Gadre et al., 2024; Hu et al., 2023; Owen, 2024; Ye et al., 2023) have proposed pipelines to estimate task-specific scaling law, while they usually incorporate models past the emergence threshold into the training set to fit a Sigmoid function and do not provide an explainable prediction of emergence abilities.
(a) Accuracy vs. effective model size (M).
(b) Binary Brier Score vs. effective model size (M).
(c) U-Shaped and inverted-U scaling.
@@
Figure 1: The accuracy, binary Brier Score, U-shaped and inverted-U scaling on the MMLU benchmark (Hendrycks et al., 2021).
@@
This paper contributes to both fronts of the literature’s discussion on emergent abilities, especially for multiple-choice tasks. First, we propose a novel procedure to describe LLM’s performance separately for each question’s difficulty level. Fig. 1 shows the result on the MMLU benchmark, with all its 14042 questions clustered into 10 groups based on their difficulties ($0\_1404\_brier$ is the group of easiest questions, and $12637\_14042\_brier$ is the hardest ones) measured by the conditionalized version of binary brier score (our adopted continuous metric discussed in Sec. 2.1.2). Performance on hard questions exhibits U-shaped scaling (Wei et al., 2023; McKenzie et al., 2023), where it worsens with scale at first, then reverses to improve with scale. In contrast, performance on easy questions exhibits an inverted U-shape followed by steady improvement with scale. The scaling trend on easy groups is consistent with the classical deep double descent of testing loss (Nakkiran et al., 2021). Moreover, the point at which performance reverts from inverse scaling to standard scaling roughly coincides with the emergence threshold beyond which model performance by traditional metrics begins to soar. Our observation could explain why LLM’s performance on many multiple-choice tasks seems stagnant for models below the emergence threshold: the scaling trend on easy questions offsets that on hard questions.
@@
This section documents LLM’s scaling trend by question difficulty level. Sec. 2.1 defines terminologies such as effective model size, emergence threshold, and our performance metrics. Sec. 2.2 describes how we group questions by difficulty level. Sec. 2.3 presents and discusses the results of the MMLU, arithmetic, and Persian-QA, which are all multiple-choice tasks.
@@ 2.1 Terminology
2.1.1 Effective Model Size and Emergence Threshold
In this paper, we refer to the effective model size $M$ as:
@@ 2.1.1 Effective Model Size and Emergence Threshold
where $C\approx 6ND$ (Kaplan et al., 2020) is the training compute, $N$ is the number of training tokens and $D$ is the number of model parameters. That is, we normalized training FLOPs $C$ to $M$ for clearer visualization. In addition, we recognize the emergence threshold $T$ as the effective model size $M$ at which the model accuracy exhibits a sharp improvement, as illustrated in Fig. 1(a).
@@ 2.1.2 Continuous Performance Metrics
where $K$ is the number of samples and $C$ is the number of classes. $p_{t,i}$ is 1 if the $t$-th sample belongs to class $i$, otherwise 0. $\hat{p}_{t,i}$ is the model’s predicted probability of the $t$-th sample being class $i$.
However, the Brier Score depends not only on the model’s predicted probability of the target choice but also on the model’s predicted probability distribution on all other choices. Since there is no a priori reason which type of distribution on non-target choices signify better ability [^1] In fact, the relation of Brier Score with accuracy measure could be noisy due to insufficient confidence calibrations (Hendrycks et al., 2021; Kapoor et al., 2024)., we propose the binary Brier Score that depends only on the predicted probability of the target choice and has an opposite sign as (Eq. 2) so that higher score means higher performance:
$Binary\_Brier=-\frac{1}{K}\sum_{t=1}^{K}\left(\hat{p}_{t,c}-1\right)^{2},$ (3)
where $\hat{p}_{t,c}$ is the model’s output probability on $t$-th sample’s correct choice $c$. For a two-class categorization where the correct choice is one class, and all incorrect choices form the other, Eq. 3 is just the standard Brier Score with the sign flipped.
We further modify the definition by using the model’s predicted probability conditional on available choices, i.e., replacing $\hat{p}_{t,c}$ in Eq. 3 with
$\frac{\hat{p}_{t,c}}{\displaystyle\sum_{c^{\prime}\in\text{ available choices}}\hat{p}_{t,c^{\prime}}}.$ (4)
This modified definition measures the model’s confidence in the correct choice relative to the sum of all available choices. Unless otherwise stated, we use this definition of binary Brier Score throughout the paper. We discuss the effect of conditionality in App. B.
(a) Accuracy vs. effective model size (M).
(b) Binary Brier Score vs. effective model size (M).
@@
Figure 2: The accuracy, binary Brier Score, U-Shaped and inverted-U scaling on the Persian-QA dataset in BIG-bench (Srivastava et al., 2023).
@@ 2.2.1 Measuring Question Difficulty Level
For a sample question $q$ of a downstream task, we define its difficulty level $D_{q}$ to be the average performance on $q$ based on the chosen performance metric (binary Brier Score in this paper) across all $L$ LLM models smaller than the emergence threshold $T$ for that downstream task:
$D_{q}=\frac{1}{L}\sum_{i=1}^{L}Binary\_Brier^{q}_{i},$ (5)
where $Binary\_Brier^{q}_{i}$ is the binary Brier Score of $i$-th LLM on sample question $q$, defined by Eq. 3.
@@ 2.3 U-Shaped and Inverted-U Scaling
(a) Accuracy vs. effective model size (M).
(b) Binary Brier Score vs. effective model size (M).
@@
Figure 3: The accuracy, binary Brier Score, U-shaped and inverted-U scaling on the arithmetic dataset in BIG-bench (Srivastava et al., 2023).
Fig. 1(a)3(a) show the scaling trend of accuracy on the MMLU, Persian-QA, and arithmetic datasets, where clear emergent phenomena are demonstrated. In contrast, Fig. 1(b)3(b) show the scaling trend when performance is measured by binary Brier Score instead. Though the scaling trend on the Persian-QA dataset is smoother, scaling trend of the binary Brier Score still exhibits a sharp increase past the emergence threshold on the MMLU and arithmetic datasets.
Fig. 1(c)3(c) show the scaling trend of the binary Brier Score with group number $G=10$. More implementation details are in App. A. Model performance on easier questions, such as $0\_1404\_brier$ in Fig. 1(c) and $0\_105\_brier$ in Fig. 2(c), displays an inverted U-shape, followed by steady improvement, i.e. performance first increases with scale, then worsens with scale, followed by a second ascent. This is consistent with the classical trend of deep double descent [^2] The term “descent” in the original paper refers to the testing loss. Hence, for the sense of accuracy or Brier Score, it is an “ascent” of performance. on testing loss (Nakkiran et al., 2021). Moreover, the reversion from inverse scaling to standard scaling roughly coincides with the emergence threshold $T$. On the contrary, performance on hard questions, such as $945\_1050\_brier$ in Fig. 2(c) and $13520\_15023\_brier$ in Fig. 3(c), displays a U-shaped scaling trend (Wei et al., 2023; McKenzie et al., 2023): model performance first decreases with scale and then consistently increases with scale. Though the scaling trend of the hardest group of questions in Persian-QA ($945\_1050\_brier$) is flat after it stops decreasing with scale, as shown in Fig. 3(c), it still starts to increase before the emergent threshold $T$.
Overall, the scaling trend of a group transitions from that of the easiest group (inverted-U followed by steady ascent) to that of the hardest group (U-shape) as we move from the easiest group to the hardest group. Because the initial scaling trends of easy questions and hard questions roughly offset each other (inverted-U vs. U-shaped), when aggregated across all difficulty levels, performance would be stagnant till the scaling trend on easy questions reverts from inverse scaling to standard scaling, followed by a sharp improvement with scale when performance on easy and hard questions both improve with scale. This could explain the emergent ability phenomenon reported in previous literature (Wei et al., 2022; Schaeffer et al., 2024a; Hu et al., 2023). We provide more results on three additional emergent tasks in App. C, scaling trend on three non-emergent tasks in App. D, and scaling trend with accuracy used as difficulty and performance measure in App. E.
@@ 3 Possible Explanation for U-shaped and Inverted-U Scaling
We provide a possible explanation for the initially opposing scaling trends (inverted-U vs. U-shaped) on easy vs. hard questions using the AI community’s previous findings (Nakkiran et al., 2021; Wei et al., 2023; McKenzie et al., 2023) in deep neural networks (DNNs)’ and specific LLMs’ behaviors.
@@ 3.1 Scaling Trend of Easy Question Groups
Figure 4: Illustration of deep double descent (Nakkiran et al., 2021) on easy question groups and U-shaped scaling (Wei et al., 2023) on hard question groups under the binary Brier Score.
As discussed in Sec. 2 and shown in Fig. 13, for a downstream task with emergent abilities, model performance on easy question groups first increases with scale, then decreases with scale, and finally reverts to increasing with scale. Fig. 4(a) illustrates the scaling trend if we flip the sign on the binary Brier Score so that a higher number means higher prediction loss. The pattern is then consistent with the deep double descent phenomenon identified in Nakkiran et al. (2021). In the context of testing error scaling law, Nakkiran et al. (2021) argues that initially, the bias-variance trade-off in the classical statistical learning theory (Hastie et al., 2009) applies, which forms the “classical regime”: complex models suffer from “overfitting” and thus, once complexity exceeds a certain threshold, models become over-sensitive to sample noises, and the effect from such bigger variance dominates the effect of further reducing testing error. On the other hand, once the model is big enough (the “modern regime”), further increase in complexity allows the model to pick from more and more interpolating models that all fit the dataset, thereby improving performance and reducing testing error to near zero.
@@ 3.2 Scaling Trend on Hard Question Group
Table 1: Examples of an easy and hard question in the MMLU benchmark. The Avg. Prob. is the average output probabilities before re-distribution over all models with effective model size $M<1.5$. Answer choices are underlined. In the hard question, small models overlook the negation “doesn’t”, giving choice C a high confidence, yet correct choice D a low confidence.
Question Description Question Group Choices Avg. Prob.
(conceptual physics, id = 44) The second law of thermodynamics tells us that heat doesn’t flow from $12637\_14042\_brier$ (hardest group) A. hot to cold ever B. cold to hot ever C. hot to cold without external energy D. cold to hot without external energy A. 0.24 B. 0.29 C. 0.29 D. 0.18
(global facts, id = 66) In 1935 roughly how many Americans were in favor of Social Security act? $0\_1404\_brier$ (easiest group) A. 90% B. 70% C. 50% D. 30% A. 0.44 B. 0.30 C. 0.17 D. 0.09
In contrast to the easy question groups, performance in hard question groups exhibits U-shaped scaling. McKenzie et al. (2023); Wei et al. (2023) have identified U-shaped scaling of LLM performance in some downstream tasks, as illustrated in Fig. 4(b). Wei et al. (2023) provides a potential explanation for the initial inverse scaling: these tasks might contain a “distractor task” that attracts models to learn to solve at first, and thus larger models perform worse. One such example is the NeQA task (McKenzie et al., 2023) , which negates each multiple-choice question in the OpenBookQA dataset (Mihaylov et al., 2018) to examine whether models would be misled by the negation. It turns out that model performance would first decline from random guesses because of the attempt to answer the non-negation part of the question. Table. 1 shows such a question in our hard question group in the MMLU benchmark. For the question “The second law of thermodynamics tells us that heat doesn’t flow from”, small models (effective model size $M<1.5$) on average assign high confidence to choice C and lowest confidence to the correct choice D, and the former is the answer if removing negation “doesn’t” from the original question.

### v2

@@ Abstract
Large language models (LLMs) have been shown to exhibit emergent abilities in some downstream tasks, where model performance stagnates at first and then improves sharply and unpredictably with scale beyond a threshold. In this work, we investigate the phenomenon by grouping questions based on difficulty level and provide a possible explanation for emergent abilities. Specifically, we observe U-shaped scaling for hard questions and inverted-U scaling followed by steady improvement for easy questions. The two scaling patterns initially offset each other, causing stagnant overall performance. The performance starts to soar when the scaling pattern of easy questions reverts from inverse to standard scaling, leading to emergent abilities. Based on this finding, we propose a simple yet effective pipeline, called Slice-and-Sandwich, to predict the emergence threshold and model performance beyond the threshold. Our code is publicly available at https://github.com/tony10101105/ExpEmergence.
@@ 1 Introduction
Large language models (LLMs) (Team et al., 2023; Achiam et al., 2023; Brown, 2020; Touvron et al., 2023a; b; Workshop et al., 2022; Li et al., 2023; Jiang et al., 2024) have shown strong potential in various downstream applications (Jumper et al., 2021; Fawzi et al., 2022; Naveed et al., 2023; Kaddour et al., 2023). Though the training-loss scaling law has been well established (Kaplan et al., 2020; Hoffmann et al., 2022), the literature is inconclusive regarding how performance on downstream tasks scales. In particular, for certain downstream tasks (Srivastava et al., 2023; Lin et al., 2022a; Pilehvar & Camacho-Collados, 2019), LLMs display emergent abilities: performance stagnates even when model training compute scales up hundredfold, and then improves sharply at an unpredictable critical threshold (Wei et al., 2022; Schaeffer et al., 2024a).
@@
Some prior work (Schaeffer et al., 2024a; b; Lu et al., 2024) link emergent abilities to crude performance metrics that fail to capture small model improvements. Hu et al. (2023) introduces the PASSUNTIL metric, showing gradual model improvement with scale. Schaeffer et al. (2024a) finds that emergent abilities mainly happen on string-match and multiple-choice tasks (Schaeffer et al., 2024a), for which traditional performance measures exhibit strong discontinuity. They propose continuous metrics such as Brier Score (Brier, 1950) and linear metrics such as token edit distance (TED) (Schaeffer et al., 2024a) to better predict LLM scaling law on downstream tasks. Schaeffer et al. (2024b) further ranks several performance metrics in correlation with the model scale. On the other hand, Michaud et al. (2024) establish the quantization model of neural scaling to explain the emergent drop of cross-entropy loss from the aspect of next-token prediction.
@@
Another focus of prior literature is the predictability of emergent abilities measured in traditional metrics like accuracy, which is crucial for monitoring and forecasting LLMs’ potentially harmful task-wise capabilities, such as writing certain malicious code (Charan et al., 2023). Though Wei et al. (2022) characterizes emergent abilities as unpredictable performance soar, some studies (Ruan et al., 2024; Gadre et al., 2024; Hu et al., 2023; Owen, 2024; Ye et al., 2023) have proposed pipelines to estimate task-specific scaling law. However, they usually incorporate models past the emergence threshold into the training set to fit a Sigmoid function and do not predict emergence abilities.
(a) Accuracy vs. log compute (M).
(b) TC Brier Score vs. log compute (M).
(c) U-Shaped and inverted-U scaling with MMLU’s questions clustered into 10 groups. Higher levels are harder questions.
@@
Figure 1: The accuracy, Target-Conditioned (TC) Brier Score, U-shaped and inverted-U scaling on the MMLU benchmark (Hendrycks et al., 2021) evaluated using 56 LLMs. Sec. 2.1.2 provides details on the TC Brier Score, which captures granular changes in model performance. App. A provides details of the 56 LLMs.
@@
This paper contributes to both fronts of discussion on emergent abilities, especially for multiple-choice tasks. First, we propose a novel procedure to evaluate the performance of LLMs on questions grouped by different difficulty levels. Fig. 1 shows the evaluation result of 56 LLMs with diverse training compute on the MMLU benchmark, whose 14042 questions are clustered into 10 groups based on their difficulty levels, with higher levels denoting harder questions. Both model performance and difficulty of each question are measured and calculated by the Target-Conditioned (TC) Brier Score (see Sec. 2.1.2 for details), which is our proposed continuous metric that is highly correlated with accuracy but can capture more granular changes in model predictions. We observe that performance on hard questions exhibits U-shaped scaling (Wei et al., 2023; McKenzie et al., 2023), where it worsens with scale at first and then reverses to improve with scale. In contrast, performance on easy questions exhibits an inverted U-shape followed by steady improvement with scale, consistent with the previously reported deep double descent of testing loss (Nakkiran et al., 2021). Moreover, the point at which performance reverts from inverse to standard scaling roughly coincides with the emergence threshold beyond which model performance begins to soar. Our observation could explain why LLM’s performance on some multiple-choice tasks stagnates for models below the emergence threshold: the scaling trend on easy questions offsets that on hard questions.
@@
This section documents LLM’s scaling trend by question difficulty level. Sec. 2.1 defines terminologies such as log compute, emergence threshold, and our performance metrics. Sec. 2.2 describes how we group questions by difficulty level. Sec. 2.3 presents and discusses the results of 6 iconic multiple-choice tasks with emergent abilities.
@@ 2.1 Terminology
2.1.1 Log Compute and Emergence Threshold
For clearer visualization, in this paper, we refer to an LLM’s log compute $M$ as:
@@ 2.1.1 Log Compute and Emergence Threshold
where $C\approx 6ND$ (Kaplan et al., 2020) is the total training compute (FLOPs) of an LLM, $N$ is the number of model parameters, and $D$ is the number of training tokens. The emergence threshold $T$ is identified manually as the log compute where the model accuracy exhibits a sharp improvement, as illustrated in Fig. 1(a).
@@ 2.1.2 Continuous Performance Metrics
where $K$ is the number of samples and $C$ is the number of classes. $p_{t,i}$ is 1 if the $t$-th sample belongs to class $i$, and 0 otherwise. $\hat{p}_{t,i}$ is the model’s predicted probability of the $t$-th sample being of class $i$.
However, the Brier Score depends not only on the model’s predicted probability of the target class (choice) but also on the predicted probability distribution of all classes. Since there is no a priori reason which type of distribution on non-target classes signify better ability  [^1] Schaeffer et al. (2024b) found that $\hat{p}^{con}_{t,c}$ (denoted as $p_{\theta}^{Choices}(\text{correct choice})$ in Schaeffer et al. (2024b))) has higher correlation with log compute than Brier Score does., we propose the Target-Conditioned (TC) Brier Score that lumps all non-target available classes into one class, conditioned on available classes, and has an opposite sign to Eq. 2 so that higher score means higher performance:
$TC\_Brier=-\frac{1}{K}\sum_{t=1}^{K}\left(\hat{p}^{con}_{t,c}-1\right)^{2},$ (3)
where $\hat{p}^{con}_{t,c}$ is the model’s predicted probability of the $t$-th sample being of target class $c$ conditional on available classes, i.e.,
$\hat{p}^{con}_{t,c}=\frac{\hat{p}_{t,c}}{\displaystyle\sum_{i\in\text{available classes}}\hat{p}_{t,i}},$ (4)
where c is the target class, $\hat{p}_{t,c}$ is the model’s predicted probability of the $t$-th sample being of the target class. We discuss the effect of conditioning in App. B.
(a) Accuracy vs. log compute (M).
(b) TC Brier Score vs. log compute (M).
@@
Figure 2: The accuracy, TC Brier Score, U-Shaped and inverted-U scaling on the Persian-QA dataset in BIG-bench (Srivastava et al., 2023).
@@ 2.2.1 Measuring Question Difficulty Level
For a question $q$ of a downstream task with emergence threshold $T$, we define its difficulty level $D_{q}$ to be the question-level TC Brier score that takes as samples the outputs on question $q$ from all $L$ LLMs smaller than the emergence threshold $T$. More specifically, we define
$D_{q}=-\frac{1}{L}\sum_{t=1}^{L}\left(\hat{p}^{con}_{t,c}-1\right)^{2}.$
@@ 2.3 U-Shaped and Inverted-U Scaling
(a) Accuracy vs. log compute (M).
(b) TC Brier Score vs. log compute (M).
@@
Figure 3: The accuracy, TC Brier Score, U-shaped and inverted-U scaling on the arithmetic dataset in BIG-bench (Srivastava et al., 2023).
Fig. 1(a)3(a) show the scaling trend of accuracy on the MMLU, Persian-QA, and arithmetic datasets, with clear ability emergence demonstrated. Fig. 1(b)3(b) show the scaling trend of TC Brier Score. Though Persian-QA has a smooth scaling, MMLU and arithmetic still exhibit a sharp increase past the emergence threshold $T$. Fig. 1(c)3(c) show the TC Brier Score scaling trend with group number $G=10$. Implementation details and model details are in App. A. Model performance on easier questions, such as difficulty level 1 in Fig. 1(c) and Fig. 2(c), displays an inverted-U shape followed by steady improvement, i.e., performance first increases and then worsens with scale, followed by a second ascent, aligning with the previously reported deep double descent [^2] The term “descent” in the original paper refers to the testing loss. Hence, for the sense of accuracy or Brier Score, it is an “ascent” of performance. on testing loss (Nakkiran et al., 2021). Moreover, the reversion from inverse scaling to standard scaling roughly coincides with $T$. In contrast, performance on hard questions, such as difficulty level 10 in Fig. 2(c) and Fig. 3(c), displays a U-shaped scaling trend (Wei et al., 2023; McKenzie et al., 2023): model performance decreases with scale in early stage and increases with scale when $M$ gets larger. Besides MMLU, arithmetic, and Persian-QA, Fig. 4 shows U-shaped vs. inverted-U scaling of Hindu knowledge, conceptual combinations, and analogical similarity datasets in Big-Bench (Srivastava et al., 2023), totaling 6 datasets, with $G=3$. Detailed results of the 3 datasets are in App. C.
Overall, the scaling trend of a question group transitions from that of the easiest group (inverted-U followed by steady ascent) to that of the hardest group (U-shape). As the initial scaling trends of easy questions and hard questions roughly offset each other when aggregated across all difficulty levels, performance stagnates until the scaling trend on easy questions reverts from inverse to standard scaling, followed by a sharp improvement when performance on easy and hard questions both improve with scale. This could explain the emergent ability phenomenon reported in previous literature (Wei et al., 2022; Schaeffer et al., 2024a; Hu et al., 2023). More results of scaling trend on 3 non-emergent tasks in App. D, and U-shaped vs. inverted-U scaling measured by accuracy are in App. F.
(a) MMLU.
(b) Arithmetic.
(c) Persian-QA.
(d) Hindu knowledge.
(e) Conceptual combinations.
(f) Analogical similarity.
Figure 4: U-shaped and inverted-U scaling on 6 datasets exhibiting emergent abilities, with group number $G=3$. Except for MMLU, the other 5 tasks are from Big-Bench. Different levels of U-shaped and inverted-U scaling trends are demonstrated across the 6 tasks.
@@ 3 Possible Explanation for U-shaped and Inverted-U Scaling
We provide a possible explanation for the initially opposing scaling trends (inverted-U vs. U-shaped) on easy vs. hard questions using the AI community’s previous findings (Nakkiran et al., 2021; Wei et al., 2023; McKenzie et al., 2023) in deep neural network and LLM behaviors.
@@ 3.1 Scaling Trend of Easy Question Groups
Figure 5: Illustration of deep double descent (Nakkiran et al., 2021) on easy question groups and U-shaped scaling (Wei et al., 2023) on hard question groups under the TC Brier Score.
As discussed in Sec. 2, for a downstream task with emergent abilities, model performance on easy questions displays an inverted-U shape followed by steady improvement. Fig. 5(a) illustrates the scaling trend if flipping the sign on the TC Brier Score so that a higher number means a higher prediction loss. The pattern is consistent with the phenomenon deep double descent identified in Nakkiran et al. (2021). In the context of testing error scaling law, Nakkiran et al. (2021) argues that initially, the bias-variance trade-off in classical statistical learning theory (Hastie et al., 2009) applies, which forms the “classical regime”: complex models suffer from “overfitting” and thus, once complexity exceeds a certain threshold, models become over-sensitive to sample noises, and the effect from such larger variance dominates the effect of further reducing testing error. On the other hand, once the model is large enough (the “modern regime”), a further increase in complexity allows the model to pick from more and more interpolating models that all fit the dataset, thus improving performance and reducing testing error to near zero.
@@ 3.2 Scaling Trend on Hard Question Group
Table 1: Examples of an easy and hard question in the MMLU benchmark. The Avg. Prob. is the average output probabilities before re-distribution over all models with log compute $M<1.5$. Answer choices (classes) are underlined. In the hard question, small models overlook the negation “doesn’t”, giving choice C a high confidence, yet correct choice D a low confidence.
Question Description Difficulty Level Choices Avg. Prob.
(conceptual physics, id = 44) The second law of thermodynamics tells us that heat doesn’t flow from level 10 (hardest group) A. hot to cold ever B. cold to hot ever C. hot to cold without external energy D. cold to hot without external energy A. 0.24 B. 0.29 C. 0.29 D. 0.18
(global facts, id = 66) In 1935 roughly how many Americans were in favor of Social Security act? level 1 (easiest group) A. 90% B. 70% C. 50% D. 30% A. 0.44 B. 0.30 C. 0.17 D. 0.09
McKenzie et al. (2023); Wei et al. (2023) have identified the U-shaped scaling of LLM in some downstream tasks, as illustrated in Fig. 5(b). Wei et al. (2023) provides a potential explanation for the initial inverse scaling: the task might contain a “distractor task” that attracts models to learn to solve at first; thus, larger models perform worse. An example is the NeQA task (McKenzie et al., 2023) , which negates each multiple-choice question in the OpenBookQA dataset (Mihaylov et al., 2018) to examine whether models would be misled by the negation. It turns out that model performance first declines from random guesses due to the attempt to answer the non-negation part of the question. Table 1 shows such a question in our hard question group in the MMLU. For the question “The second law of thermodynamics tells us that heat doesn’t flow from”, small models ($M<1.5$) on average assign high confidence to choice C and lowest confidence to correct choice D, where the former is the answer if removing negation “doesn’t” from the original question. More MMLU questions to demonstrate U-shaped and inverted-U scaling trends w.r.t. model log compute are in App. E.